# Infinity-Parser: Layout-Aware Reinforcement Learning for Scanned Document Parsing

### v1

@@ Abstract
Automated parsing of scanned documents into richly structured, machine-readable formats remains a critical bottleneck in Document AI, as traditional multi-stage pipelines suffer from error propagation and limited adaptability to diverse layouts. We introduce layoutRL, an end-to-end reinforcement-learning framework that trains models to be explicitly layout-aware by optimizing a composite reward of normalized edit-distance, paragraph-count accuracy, and reading-order preservation. Leveraging our newly released dataset—Infinity-Doc-55K, which combines 55K high-fidelity synthetic scanned document parsing data with expert-filtered real-world documents—we instantiate layoutRL in a vision-language-model–based parser called Infinity-Parser. Evaluated on English and Chinese benchmarks for OCR, table and formula extraction, and reading-order detection, Infinity-Parser achieves new state-of-the-art performance in both accuracy and structural fidelity, outpacing specialist pipelines and general-purpose vision-language models. We will publicly release our code and dataset to accelerate progress in robust document understanding.
@@ 1 Introduction
Automating the parsing of scanned documents into layout, machine-readable formats is a longstanding challenge in Document AI [19, 52, 59, 62, 69]. While scanned documents provide fine-grained control over visual layouts, they lack explicit semantic markers for essential elements such as paragraphs, headings, tables, and lists. This absence introduces structural ambiguity, making it difficult to recover the hierarchical organization of documents and limiting the scalability of traditional rule-based or multi-stage parsing systems [52, 37]. Conventional approaches typically rely on a sequence of supervised tasks—including layout detection, optical character recognition (OCR), table recognition and math formula recognition —followed by heuristic or scripted post-processing to reconstruct document layout [6, 61, 25, 59, 2, 8]. These pipelines not only demand extensive annotated resources but also suffer from error propagation and limited adaptability to diverse or unseen layouts.
Recent advances in vision-language models (VLMs) have enabled end-to-end approaches that directly map visual document inputs to layout outputs such as Markdown [3, 8, 63, 48]. However, off-the-shelf pretraining objectives—focused on image–text alignment or language modeling—do not inherently encourage preserving a document’s spatial arrangement or reading order. These approaches simplify traditional multi-stage pipelines and offer new possibilities for document parsing, existing models still struggle to generalize effectively on downstream tasks [52, 59]. A key reason lies in the fundamental mismatch between pretraining objectives and the structural understanding required for document parsing—most pretraining focuses on image-text alignment and reading comprehension, without optimizing for the structural complexities inherent in documents. Compared to standard text, documents exhibit richer, more intricate layouts, characterized by hierarchical layouts, nested elements, and multimodal content [17, 44, 65, 64]. This complexity introduces challenges that conventional token-level supervision cannot address: prediction errors on a few critical tokens can propagate into substantial structural deviations, and current methods lack fine-grained, layout-aware signals to guide models toward outputs that are both semantically accurate and structurally faithful [11, 24, 42]. Recent studies show that reinforcement learning (RL) achieves stronger generalization in rule-based and visual reasoning tasks [9]. Existing RL approaches remain constrained by coarse-grained, binary outcome rewards that fail to provide the fine-grained, layout-aware supervision needed for modeling complex document layouts [29, 56, 73, 32]. Designing effective layout-aware reward functions tailored for document parsing remains an underexplored challenge.
Figure 1: Overview of the layout-aware reinforcement learning framework. It includes data preparation, reinforcement finetuning, and reward calculation based on edit distance, count, and order to enhance structural understanding.
@@
To address these limitations, we propose layoutRL, the first end-to-end reinforcement-learning framework that makes document parsers explicitly layout-aware, and construct a new dataset, as shown in Figure 1. Specially, first, to overcome the generalization gap caused by the mismatch between pretraining objectives and structural understanding, we incorporate verifiable rewards [43, 24, 42, 66] that provide document-level supervision beyond token-level signals, encouraging the model to learn transferable structural representations. Second, to mitigate structural hallucination under complex layouts, we design Edit Distance Reward [21] and Layout Parsing Reward signals, including count and order that enforce fine-grained alignment between predictions and ground-truth layouts, improving layout coherence. Additionally, to train our RL model we construct Infinity-Doc-55K, a 55 K-document corpus providing large-scale, high-quality supervision. It combines (1) high-fidelity synthetic scanned document parsing data—generated via HTML templates and browser rendering—and (2) expert-filtered real-world samples, pseudo-labeled through a cross-model agreement pipeline to capture genuine layout diversity. Finally, we train a vision-language-model–based parser, Infinity-Parser, achieves state-of-the-art results across English and Chinese benchmarks—covering OCR, table and formula extraction, and reading-order detection—and consistently outperforms both specialist pipelines and leading vision-language models.
@@
  • We propose layoutRL, a new reinforcement learning framework for end-to-end scanned document parsing, which explicitly trains models to be layout-aware by optimizing verifiable, multi-aspect rewards. Our multi-aspect reward design combines normalized edit distance, paragraph count accuracy, and reading order preservation, enabling stable training and improved structural robustness.
  • We introduce Infinity-Doc-55K, a large-scale dataset of 55,066 scanned documents that combines high-quality synthetic data with diverse real-world samples. The dataset features rich layout variations and comprehensive structural annotations, enabling robust training of document parsing models.
@@
Benchmark Document Domain Annotation Type End-to-End Task Exactly Match
@@
Fox [25] 2
Nougat [6] 1
GOT OCR 2.0 [59] 2
OmniDocBench [36] 9
@@
DocStruct4M [13] -
olmOCR-mix [39] -
Infinity-Doc-55K 7
Table 1: A comparison between Infinity-Doc-55K and existing datasets. BBox: Bounding boxes. Text: Text in Unicode. Table: Table in LaTeX/HTML/Markdown. Formula: Formula in LaTeX. Attributes: Page- and BBox-Level Attributes. OCR: Optical Character Recognition; TR: Table Recognition; MFR: Math Formula Recognition; ROD: Reading Order Detection. Multi-Type Doc: Whether the dataset includes documents from multiple domains or categories.
@@ 2 Methodology
In this section, we first introduce Infinity-Doc-55K, our large-scale multimodal dataset for end-to-end scanned document parsing. We then describe our rule-based multi-aspect reward framework, which integrates edit distance, paragraph count, and order criteria under a unified reinforcement learning objective optimized via Group Relative Policy Optimization (GRPO).
2.1 Infinity-Doc-55K and Generation Pipelines
@@
We introduce Infinity-Doc-55K, a large-scale, multimodal dataset of 55,066 richly annotated documents for end-to-end scanned document parsing. Unlike prior benchmarks that target isolated subtasks (e.g., layout detection, OCR, or table recognition), Infinity-Doc-55K provides holistic supervision by pairing rendered scanned documents pages with their ground-truth Markdown representations. This design enables training and evaluating models that directly translate visual inputs to layout outputs without relying on brittle, multi-stage pipelines. As shown in Table 1, compared to existing works, Infinity-Doc-55K not only significantly enhances task diversity but also substantially improves overall data quality through our proposed synthetic generation mechanism. More details on the data distribution are provided in the Data Details section of the Appendix.
Figure 2: Data construction pipelines for document parsing. (a) Real-world pipeline enhance quality by combining multiple expert models and layout analysis, yielding better-aligned supervision through intersection and reading order reasoning. (b) Synthetic pipeline leverages structured HTML templates and browser rendering to generate clean, exactly-aligned scanned document parsing data, ensuring high-quality supervision for end-to-end parsing.
@@
To construct Infinity-Doc-55K, we design a dual-pipeline framework that integrates both synthetic and real-world document generation, as illustrated in Figure 2. This design addresses a critical limitation of traditional data construction pipelines, which often rely on weak supervision and pseudo-labeling from a single model applied to crawled scanned documents. These pipelines frequently suffer from noisy, misaligned, or incomplete annotations, especially in complex layouts or multilingual content, thus hindering model performance and generalization. To overcome these issues, our dual-pipeline framework is motivated by the need to balance annotation quality and structural diversity. The synthetic branch provides highly accurate, clean, and precisely aligned annotations at scale, while the real-world branch introduces naturally occurring layout variability and semantic richness, which are essential for building models that generalize robustly in practical applications.
@@
Real-World Data We develop a real-world data construction pipeline to capture the structural complexity and natural layout variability of documents across practical domains. We collect diverse scanned documents from sources such as financial reports, medical records, academic papers, books, magazines, and web pages, covering both dense and sparse content layouts. To generate structural annotations, we adopt a multi-expert strategy, where specialized models handle different structural elements, such as layout blocks, texts, formulas, and tables. For example, overall layouts are analyzed by visual layout model [18], formula regions are processed by dedicated formula recognition model [50], tables are parsed by transformer-based table extractor [6]. A cross-validation mechanism is then applied to filter out inconsistencies by comparing the outputs of expert models and VLM [1]. Only regions with consistent predictions across models are retained as high-confidence pseudo-ground-truth annotations. This layout-aware filtering results in a rich and reliable dataset that reflects the complexity of real-world documents and supports robust document parsing model training.
@@
Synthetic Data We design a synthetic data construction pipeline. We collect text and images from sources such as Wikipedia, web crawlers, and online corpora, and use Jinja [34] templates to inject sampled content into predefined single-, double-, or triple-column HTML layouts. These pages are rendered into scanned documents using a browser engine, followed by automated filtering to remove low-quality or overlapping images. Ground-truth annotations are extracted by parsing the original HTML to produce aligned Markdown representations. This synthetic approach not only significantly reduces construction costs and ensures annotation accuracy and structural diversity, but more importantly, it addresses the longstanding issue of imprecise or inconsistent supervision commonly found in pseudo-labeled datasets, providing high-quality and well-aligned supervision for training end-to-end models.
Data Details As illustrated in Figure 3, the dataset spans seven diverse document domains, making it one of the most richly annotated and structurally varied resources to date. Each domain is represented by two sample pages, highlighting the broad variability in layout design, content structure, and semantic density. For instance, Medical Reports typically contain structured tables with clinical measurements and diagnostic notes. Synthetic Documents are algorithmically generated to replicate real-world formats, providing layout diversity for training robust parsers. Financial Reports feature dense tables and formal accounting records, while Academic Papers often follow two-column layouts with references, equations, and figures. Books combine narrative content with visual illustrations, and Magazines blend images and stylized text for reader engagement. Finally, Web Pages, when saved as PDFs, preserve HTML-based structures that integrate tables, lists, and dynamic elements. This visual taxonomy exemplifies the structural and semantic diversity present in real-world documents, highlighting the core challenge faced by document AI systems: reliably parsing heterogeneous layouts and extracting structured information across a wide variety of formats.
Figure 4: Overview of the Infinity-Parser training framework. Our model is optimized via reinforcement fine-tuning with edit distance, layout, and order-based rewards. The detailed formulations of these reward functions are provided in Section 2.2.
Table 13 provides an overview of the document types included in the Infinity-Doc-55K dataset, detailing the composition across both real-world and synthetic sources. The real-world portion consists of 48.6k documents spanning six domains: financial reports, medical reports, academic papers, books, magazines, and web pages. These documents were collected from the web and annotated using a pseudo-labeling pipeline based on expert model agreement. While this approach enables large-scale data acquisition, the resulting label quality is relatively low due to occasional inconsistencies across models. Additionally, real-world data collection incurs a high cost, especially in terms of manual filtering, formatting normalization, and layout validation.

### v3

@@ Abstract
Document parsing from scanned images into structured formats remains a significant challenge due to its complexly intertwined elements such as text paragraphs, figures, formulas, and tables. Existing supervised fine-tuning methods often struggle to generalize across diverse document types, leading to poor performance, particularly on out-of-distribution data. This issue is further exacerbated by the limited availability of high-quality training data for layout-aware parsing tasks. To address these challenges, we introduce LayoutRL, a reinforcement learning framework that optimizes layout understanding through composite rewards integrating normalized edit distance, paragraph count accuracy, and reading order preservation. To support this training, we construct the Infinity-Doc-400K dataset, which we use to train Infinity-Parser, a vision-language model demonstrating robust generalization across various domains. Extensive evaluations on benchmarks including OmniDocBench, olmOCR-Bench, PubTabNet, and FinTabNet show that Infinity-Parser consistently achieves state-of-the-art performance across a broad range of document types, languages, and structural complexities, substantially outperforming both specialized document parsing systems and general-purpose vision-language models. We will release our code, dataset, and model to facilitate reproducible research in document parsing.
22footnotetext: Lead contributor. Contact: zuminghuang@gmail.com33footnotetext: Corresponding Author.
@@ 1 Introduction
Document parsing aims to convert scanned documents into structured, machine-readable formats and represents one of the core tasks in document intelligence (Hwang et al., 2021; Wang et al., 2024b; Wei et al., 2024; Xia et al., 2024; Zhang et al., 2024). Unlike traditional OCR that focuses solely on text recognition, document parsing requires comprehensive recovery of hierarchical document structures, including the dependency relationships among elements such as paragraphs, headers, tables, and formulas—a capability that is crucial for downstream applications including legal contract analysis, scientific literature mining, and financial report processing. Traditional approaches typically rely on multi-stage pipelines that decompose the task into supervised sub-tasks—such as layout detection, OCR, table recognition, and formula recognition—followed by heuristic post-processing to reconstruct document structure (Blecher et al., 2024; Wei et al., 2025; Liu et al., 2024b; Wei et al., 2024; Bai et al., 2024; Chen et al., 2024). However, such pipeline-based methods are prone to error propagation and exhibit limited adaptability when confronted with diverse layout variations.
Recent approaches primarily reformulate document parsing as end-to-end perception tasks using vision-language models (VLMs) trained through supervised fine-tuning (SFT). However, this paradigm faces fundamental limitations. Although SFT provides token-level supervision, it often overfits to surface patterns rather than learning generalizable structural representations. This limitation is further compounded by the scarcity of large-scale, high-quality training data for document parsing, which hinders models from acquiring layout-aware knowledge. To overcome these limitations, reinforcement learning (RL) presents a promising alternative, having demonstrated strong generalization capabilities in vision and multimodal tasks, where outcome-based rewards help models learn transferable representations (Huang et al., 2025; Liu et al., 2025a; Wang et al., 2025c). This raises a fundamental question: can reinforcement learning drive models toward generalizable layout parsing rules? Unfortunately, current RL approaches remain limited by coarse-grained, binary outcome rewards that fail to provide the fine-grained, layout-aware supervision necessary for modeling complex document layouts (Guo et al., 2025; Shao et al., 2024). Therefore, how to effectively apply RL to document parsing remains an underexplored yet critical challenge.
Figure 1: Comparison of document parsing performance on OmniDocBench under different training strategies as training data size increases. Left: Evaluation with two complementary metrics: (1) Paragraph-level accuracy (edit distance evaluation on element contents only), which assesses element-wise consistency within individual element contents, independent of inter-element reading order; and (2) Page-level accuracy (edit distance evaluation on element contents and reading order), which measures global document reconstruction quality by aligning predicted outputs (e.g., texts, tables, and formulas) with ground-truth sequences. Right: In-Distribution and Out-of-Distribution task performance measured by accuracy score (1 – NED). See detailed descriptions of the task in Section $4.$
@@
To address the challenge of effectively applying RL algorithms to document parsing tasks and the lack of large-scale, high-quality data in industry to support this training process, we propose LayoutRL, the first end-to-end reinforcement learning framework for layout-aware document parsing, and construct Infinity-Doc-400K, a large-scale dataset. Specifically, our approach does not rely on explicit reasoning processes, but rather treats the entire document parsing result as the final answer and guides model learning through carefully designed reward mechanisms. We introduce verifiable rewards (Shen et al., 2025; Liu et al., 2024a; Shao et al., 2024; Yaowei Zheng, 2025), which consist of Edit Distance Reward (Levenshtein et al., 1966) and Layout Parsing Reward signals, enforcing fine-grained alignment between predictions and ground-truth layouts and providing document-level supervision beyond token-level signals to encourage the model to learn transferable structural representations. Additionally, we construct a 400K-document corpus providing large-scale, high-quality supervision. It combines (1) high-fidelity synthetic scanned document parsing data—generated via HTML templates and browser rendering—and (2) expert-filtered real-world samples, pseudo-labeled through a cross-model agreement pipeline to capture genuine layout diversity. Built on this dataset, we train Infinity-Parser, an end-to-end VLM-based parser that directly outputs structured document representations.
To demonstrate the effectiveness of our approach, we conduct extensive experiments across multiple document parsing benchmarks. Results in Figure 1 show that while SFT achieves strong paragraph-level accuracy, its page-level performance struggles to improve as data size increases, reflecting its tendency to memorize surface patterns rather than model the hierarchical dependencies between document structures and elements. In contrast, our layout-aware reinforcement learning method significantly outperforms SFT in page-level accuracy, while maintaining competitive paragraph-level performance. These findings demonstrate that verifiable, layout-aware rewards enable models to move beyond simple token imitation, achieving consistent improvements in both local detail fidelity and global structural understanding, thereby establishing a more robust paradigm for document parsing. Moreover, our method also exhibits strong generalization to unseen task types in downstream applications. Finally, our method, Infinity-Parser, achieves new state-of-the-art results on four diverse benchmarks—OmniDocBench, olmOCR, PubTabNet, and FinTabNet—highlighting its effectiveness and strong cross-domain generalization.
@@
  • We propose LayoutRL, a new reinforcement learning framework for end-to-end scanned document parsing, which explicitly trains models to be layout-aware by optimizing verifiable, multi-aspect rewards. Our multi-aspect reward design combines normalized edit distance, paragraph count accuracy, and reading order preservation, improving structural robustness.
  • We introduce Infinity-Doc-400K, a large-scale dataset of 400,482 scanned documents that combines high-quality synthetic data with diverse real-world samples. The dataset features rich layout variations and comprehensive structural annotations, enabling robust training.
@@
2 Related Work
2.1 Reinforcement Learning for Language Models
Recent advancements in Large Language Models (LLMs) such as OpenAI’s GPT series (OpenAI, 2024)), DeepSeek-R1 (Guo et al., 2025), and Gemini (Team et al., 2023) have highlighted the significant potential of Reinforcement Learning (RL) in enhancing their reasoning capabilities. This RL paradigm has been successfully extended to other domains demanding sophisticated reasoning, including code generation (Li et al., 2022b; Zeng et al., 2025), autonomous tool utilization (Schick et al., 2023; Wang et al., 2025a), and information retrieval (Nakano et al., 2021). Similarly, RL has demonstrated its efficacy in the domain of Visual Language Models (VLMs), including precise object counting (Peng et al., 2025), nuanced visual perception (Liu et al., 2025b), and complex multimodal reasoning (e.g., VL-Rethinker (Wang et al., 2025b), Pixel Reasoner (Su et al., 2025), Vision-R1 (Huang et al., 2025)). These pioneering works have predominantly relied on binary outcome rewards to guide RL training. Complementary to these efforts, our work demonstrates the effectiveness of incorporating layout-aware and layout-based rewards for document parsing, offering a more granular and contextually relevant feedback mechanism.
2.2 VLM-based Document Parsing
Recent advancements in document understanding and optical character recognition (OCR) have highlighted their importance as critical benchmarks for evaluating the perceptual capabilities of vision-language models (VLMs). By incorporating large-scale OCR corpora during pretraining, models such as GPT-4o (Achiam et al., 2023) and Qwen2-VL (Bai et al., 2024) have achieved competitive performance on document content extraction tasks. Building upon these foundations, the emergence of VLMs has further accelerated the progress of end-to-end document parsing, giving rise to a range of models such as Donut (Blecher et al., 2024), Nougat (Blecher et al., 2023), Kosmos-2.5 (Lv et al., 2024), Vary (Wei et al., 2025), mPLUG-DocOwl (Hu et al., 2024b), Fox (Liu et al., 2024b), and GOT (Wei et al., 2024). These models have continued to improve their understanding of visual layouts and textual content by leveraging advancements in visual encoders (Dosovitskiy et al., 2020), language decoders (Bai et al., 2024), and data construction pipelines. Despite the success of these VLM-based approaches in enabling end-to-end document parsing, they still face generalization challenges on downstream layout parsing tasks (Wang et al., 2024b). To address this issue, we propose leveraging reinforcement learning to provide a more effective training paradigm that better aligns with the demands of document parsing.
Benchmark
Document
Domain
Annotation Type End-to-End Task Exactly Match
@@ 2.2 VLM-based Document Parsing
Fox (Liu et al., 2024b) 2
Nougat (Blecher et al., 2024) 1
GOT OCR 2.0 (Wei et al., 2024) 2
OmniDocBench (Ouyang et al., 2024) 9
@@
DocStruct4M (Hu et al., 2024a) -
olmOCR-mix (Poznanski et al., 2025) -
Infinity-Doc-400K 7
Table 1: A comparison between Infinity-Doc-400K and existing datasets. BBox: Bounding boxes. Text: Text in Unicode. Table: Table in LaTeX/HTML/Markdown. Formula: Formula in LaTeX. Attributes: Page- and BBox-Level Attributes. OCR: Optical Character Recognition; TR: Table Recognition; MFR: Math Formula Recognition; ROD: Reading Order Detection. Multi-Type Doc: Whether the dataset includes documents from multiple domains or categories.
@@ 3 Methodology
In this section, we first introduce Infinity-Doc-400K, our large-scale multimodal dataset for end-to-end scanned document parsing. We then describe our rule-based multi-aspect reward framework, which integrates edit distance, paragraph count, and order criteria under a unified reinforcement learning objective optimized via Group Relative Policy Optimization (GRPO).
3.1 Infinity-Doc-400K and Generation Pipelines
@@
We introduce Infinity-Doc-400K, a large-scale, multimodal dataset of 400,066 richly annotated documents for end-to-end scanned document parsing. Unlike prior benchmarks that target isolated subtasks (e.g., layout detection, OCR, or table recognition), Infinity-Doc-400K provides holistic supervision by pairing rendered scanned document pages with their ground-truth Markdown representations. This design enables training and evaluating models that directly translate visual inputs to layout outputs without relying on brittle, multi-stage pipelines. As shown in Table 1, compared to existing works, Infinity-Doc-400K not only significantly enhances task diversity but also substantially improves overall data quality through our proposed synthetic generation mechanism. More details on the data distribution and quality control are provided in the Data Details section of the Appendix.
@@
To construct Infinity-Doc-400K, we design a dual-pipeline framework that integrates both synthetic and real-world document generation, as illustrated in Figure 2. This design addresses a critical limitation of traditional data construction pipelines, which often rely on weak supervision and pseudo-labeling from a single model applied to crawled, scanned documents. These pipelines frequently suffer from noisy, misaligned, or incomplete annotations, especially in complex layouts or multilingual content, thus hindering model performance and generalization. To overcome these issues, our dual-pipeline framework is motivated by the need to balance annotation quality and structural diversity. The synthetic branch provides highly accurate, clean, and precisely aligned annotations at scale, while the real-world branch introduces naturally occurring layout variability and semantic richness, which are essential for building models that generalize robustly in practical applications.
Real-World Data
@@
We develop a real-world data construction pipeline to capture the structural complexity and natural layout variability of documents across practical domains. We collect diverse scanned documents from sources such as financial reports, medical records, academic papers, books, magazines, and web pages, covering both dense and sparse content layouts. To generate structural annotations, we adopt a multi-expert strategy, where specialized models handle different structural elements, such as layout blocks, texts, formulas, and tables. For example, overall layouts are analyzed by a visual layout model (Huang et al., 2022), formula regions are processed by a dedicated formula recognition model (Wang et al., 2024a), and tables are parsed by a transformer-based table extractor (Blecher et al., 2024). A cross-validation mechanism is then applied to filter out inconsistencies by comparing the outputs of expert models and VLMs. Only regions with consistent predictions across models are retained as high-confidence pseudo-ground-truth annotations. This layout-aware filtering results in a rich and reliable dataset that reflects the complexity of real-world documents and supports robust document parsing model training.
Synthetic Data
@@
We design a synthetic data construction pipeline. We collect text and images from sources such as Wikipedia, web crawlers, and online corpora, and use Jinja (Nipkow, 2003) templates to inject sampled content into predefined single-, double-, or triple-column HTML layouts. These pages are rendered into scanned documents using a browser engine, followed by automated filtering to remove low-quality or overlapping images. Ground-truth annotations are extracted by parsing the original HTML to produce aligned Markdown representations. This synthetic approach not only significantly reduces construction costs and ensures annotation accuracy and structural diversity, but more importantly, it addresses the longstanding issue of imprecise or inconsistent supervision commonly found in pseudo-labeled datasets, providing high-quality and well-aligned supervision for training end-to-end models.
Quality Control Measures
To ensure reliable annotations at scale, we designed a hybrid quality control strategy for Infinity-Doc-400K. First, three domain experts with doctoral degrees in document analysis manually inspected about 5% of the data. Their feedback not only identified potential errors but also served as a quality anchor for evaluating annotation consistency. Guided by these inspections, we iteratively refined the screening rules no fewer than five times, continuously improving the labeling pipeline. Finally, to achieve scalability, we employed a model-based cross-verification mechanism: multiple models generated annotations for real-world samples, with high-consistency outputs retained and inconsistent cases fed back into further rule refinement. This layered framework—anchored by expert inspection, strengthened through iterative rule optimization, and scaled via model cross-checking—effectively balances annotation reliability with dataset scalability.
Figure 2: Data construction pipelines for document parsing. (a) Real-world pipelines enhance quality by combining multiple expert models and layout analysis, yielding better-aligned supervision through intersection and reading order reasoning. (b) Synthetic pipeline leverages structured HTML templates and browser rendering to generate clean, exactly-aligned scanned document parsing data, ensuring high-quality supervision for end-to-end parsing.